Two insulated charged copper spheres A and B of identical size have charges qA and qB respectively.When they are brought in contact with each other and finally separated, what are the new charges on them?


Grade:12

1 Answers

Thareeq Roshan
13 years ago

Here after the contact ,both the surfaces of the sphere will have the same potential .so,

assume that q(a) > q(b)

let x be the charge lost by sphere A and which is gained by sphere B

 

K(q(a)-x)/r2 =  K(q(b)+x)/r2       

 

q(a)-x=q(b)+x

 

[q(a)-q(b)]/2=x

 

the new charge on,

SPHERE A = q(a) -[q(a)-q(b)]/2

              = [q(a)+q(b)]/2

 

SPHERE B=q(b)+[q(a)-q(b)]/2

                    =[q(a)+q(b)]/2